a spherical shell of charge of uniform density . Plot E for distance r from the center of the shell ranging from zero to 30 cm. Assume that  = 1.010-6 C/m3, a = 10 cm and b = 20 cm.
1
Expert's answer
2019-12-20T09:55:43-0500

If the shell is made of a non-conducting material, the field inside the shell (if a=10 cm is the radius of the cavity) is zero:


"0<r<a: E_1=0."

The field outside the shell is easy to find and it's described with the same equation as for a point charge:


"r>b: E_3=\frac{1}{4\pi\epsilon_0}\cdot Q/r^2,\\
\space\\
Q=\rho\cdot\frac{4}{3}\pi b^3,\\
\space\\
E_3=\frac{\rho b^3}{3\epsilon_0r^2}."

The field within the material inside b-a = 20-10 = 10 cm can be described with the equation shown below that can be derived from Gauss's law:


"a<r<b: E_2=\frac{(r^3-a^3)\rho}{3r^2\epsilon_0}."

The field: